Diffusion Morning Report 05/27/04

ted berg tberg at snf.stanford.edu
Thu May 27 06:37:30 PDT 2004

Good Morning all,
    Looking fair this morning, BPSG is back together after replacing TC 
sheaths and injectors- needs a test. Poly is waiting for a tube to change.
TEOS 2 began testing looks like recipe will need some tweeks. Nitride 
still under observation. Everything else looiking OK

Diffusion 	Date: 	5/27/2004 	 
Tube Name/Number 	Status 	Comments 	THICKNESS-Change at
Tylan 1 	Up 	  	 
Tylan 2 	Up 	  	 
Tylan 3 	Up 	  	 
Tylan 4 	Up 	  	 
Tylan 5-BBr3 	Up 	  	 
Tylan 6-POCl3 	Up 	  	 
Tylan 7-TEOS2 	down 	Qual testing to begin 	 
Tylan 8-SiGe 	Up 	  	6.6/30
Tylan 9-Poly 	down 	Scheduled Pull and Clean 	0/30
Tylan 10-Nitride 	Up 	Particles? 	12.6/30
Tylan 11-BPSG-LTO 	yellow 	 needs test 	7.5/30
Tylan 12-FGA 	Up 	  	 
Tystar 	Up 	  	12.6/30
Doped Poly 	GONE 	  	 
TEOS 	yellow 	 running 	 
Tylan 16 	N/A 	  	 
Burn Box 	Up 	  	 
ZYGO 	Up 	  	 
GAAS 22 	UP 	  	 
GAAS 23 	UP 	  	 
SiH4 for furnaces 	  	Auto changeover working well 	A on line 500 PSI

-------------- next part --------------
An HTML attachment was scrubbed...
URL: <http://snf.stanford.edu/pipermail/diffusion/attachments/20040527/28a48968/attachment.html>

More information about the diffusion mailing list